Calculate the time taken for a car to cover a distance of 100m if the initial speed is 10km/h and it constant acceleration of 1.5m/s


1
Expert's answer
2022-04-07T18:35:36-0400

"s=vt+\frac{at^2}2,"

"t=\frac{\sqrt{v^2+2as}-v}{a},"

"t=9.7~s."